A 17' by 10' rectangular garden is divided into 4 parts. 2 parts have equal areas, the 3rd is twice as large as each of the first 2, and the 4th part's area is 20 sq. ft. than each of the first 2. Determine area of each of the 4 parts.


Total area = 170 sq. ft.

4 parts are A, B, C, D.
A = B.
C = 2A, 2B.

D = A + 20, B + 20.
Not sure how to continue. Non-homework.

Found 3 solutions by Boreal, ankor@dixie-net.com, MathTherapy:
Answer by Boreal(15235)

170=A+B+C+D=A+A+2A+A+20=5A+20
5A=150
A=30
B=30
C=60
D=50
units ft^2

Answer by ankor@dixie-net.com(22740)

A 17' by 10' rectangular garden is divided into 4 parts.
2 parts have equal areas, the 3rd is twice as large as each of the first 2, and the 4th part's area is 20 sq. ft. than each of the first 2.
Determine area of each of the 4 parts.
:
Find the total area: 17 * 10 = 170 sq/ft
a^2 + a^2 + 2a^2 + 20 = 170
4a^2 = 170 - 20
4a^2 = 150
a^2 = 150/4
a^2 = 37.5
the four parts: 37.5 + 37.5 + 75 + 20 = 170

Answer by MathTherapy(10549)


A 17' by 10' rectangular garden is divided into 4 parts. 2 parts have equal areas, the 3rd is twice as large as each of the first 2, and the 4th part's area is 20 sq. ft. than each of the first 2. Determine area of each of the 4 parts.


Total area = 170 sq. ft.

4 parts are A, B, C, D.
A = B.
C = 2A, 2B.

D = A + 20, B + 20.
Not sure how to continue. Non-homework.